[Brahms-l] Shift report 20040113 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue Jan 13 2004 - 00:04:41 EST
------------------------------------------------------------
Date: Tue Jan 13 00:04:40 2004, Shift: 16:00-24:00
Supervisor: rw,or,cr, with: 
Log book#15 pages 31 - 31

Shift Summary:
------------------------------------------------------------
1.Good store from 21:30, still running

2. FS 4deg. D1=2442B   MRS 90deg. D5=2255A
    run         Trg 3     Trg 6
    9649        75k       225k  
    9650        70k       210k

3. db put the RJIC clock on the ZDC trigger just before this run but it did 
work. So the clock is taken out for the above run.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jan 13 00:04:47 2004

This archive was generated by hypermail 2.1.8 : Tue Jan 13 2004 - 00:05:06 EST